OpenStudy (anonymous):

Two children are riding on a merry-go-round. Child A sits at a greater distance from the axis of rotation than child B. Which child has the larger centripetal velocity? A) A B) B C) They have the same non-zero centripetal velocity. D) There is no way to determine without knowing the applied force. I chose B, because it has a smaller distance from the axis of rotation. I'd assume that a greater distance would decrease the number. Did I think it through wrong?

OpenStudy (anonymous):

there is nothing called centripetal velocity if there is then it is always zero in circular motion...

OpenStudy (mos1635):

wich is centripetal velocity???

OpenStudy (anonymous):

for linear velocity and centripetal acceleration,they are more for A

OpenStudy (anonymous):

Of course. I was trying to remember centripetal acceleration. But sadly one of my partners has my notes so I can't look over at those right away. Thank you so much!

OpenStudy (anonymous):

you are welcome!

OpenStudy (mos1635):

a=v^2/r a=(ω*r)^2/r a=ω^2*r r1>r2 =>a1>a2

OpenStudy (anonymous):

P.S:omega (angular velocity) is same for both..and rest is as mos163 said..

OpenStudy (mos1635):

+1472